OpenStudy (anonymous):

evaluate the integral ∫1/[16-x2(squared) )5/2].dx

OpenStudy (raspberryjam):

\[\int\frac{1}{16-2x^2}\] ? Confused about that 5/2 on the side there...

OpenStudy (anonymous):

( )5/2 that is squared of the bracket

OpenStudy (raspberryjam):

oooh like \[\int\frac{1}{{(16-2x)}^{5/2}}\] ?

OpenStudy (2bornot2b):

Evaluate the integral \[\int\frac{1}{{(16-2x)}^{5/2}}dx\]

OpenStudy (raspberryjam):

Make it simpler. Turn it into: \[\int(16-2x)^{-\frac{5}{2}}dx\] which is the same thing.

OpenStudy (anonymous):

this is the sum\[\int\limits1/(16-x^2)^{5/2}\].dx

OpenStudy (anonymous):

thanks for ur reply

OpenStudy (anonymous):

can u solve this sum?

OpenStudy (anonymous):

it's difficult to find the answer

OpenStudy (raspberryjam):

Argh it's a trigonometry integral...I really hate these types of problems too. \[\int\frac{1}{\sqrt{(16-x^2)^5}}\] is what you've got. This should look familiar to you...look at the derivatives of arcsin, arctan, arccos, etc.

OpenStudy (raspberryjam):

oops I forgot the dx in the integral..just pretend its there :P

OpenStudy (anonymous):

i too hate those problems.the sum is like this.. \[\int\limits1/(16-x ^{2})\left(\begin{matrix}5/2 \\ ?\end{matrix}\right)\].dx (no question mark below 5/2)

OpenStudy (anonymous):

it's difficult to type the sums here

OpenStudy (raspberryjam):

\[\int\frac{1}{\sqrt{(16-x^2)^5}}dx\] This is the same thing as what you have written. Look at the derivatives of arcsin, arctan, arccos, etc. One of them should look similar.

OpenStudy (raspberryjam):

\[(16-x^2)^{5/2}=\sqrt{(16-x^2)^5}\]

OpenStudy (anonymous):

ooh that's right

OpenStudy (anonymous):

but,when we solve this, \[\sqrt{16}(1-\sin ^{2}\theta)^{5}\] what is the answer? (square root of all)

OpenStudy (anonymous):

it can't be 4cos theta because root 5 is there outside the 1- sin squared theta bracket

OpenStudy (raspberryjam):

Try looking through these: http://www.analyzemath.com/calculus/Differentiation/inverse_trigonometric.html

OpenStudy (raspberryjam):

One of them should look close...

OpenStudy (anonymous):

difficult to find.

OpenStudy (raspberryjam):

try \[x=\frac{1}{4} sin x\] and \[dx=\frac{1}{4} cos dx\]

OpenStudy (anonymous):

x=4sin theta dx=4costheta.dtheta

OpenStudy (anonymous):

\[x=4\sin \theta\] \[dx=4\cos \theta.d \theta\]

OpenStudy (anonymous):

i can solve up to that ..i need how to solve the final part.at last i got\[1/4 \int\limits \cos ^{-4} \theta.d \theta\]

OpenStudy (anonymous):

then after that????

OpenStudy (raspberryjam):

Hmm well actually if you make x=4 sin x and dx=4 cos dx it should work out ok because then you get: \[\int\frac{4 cos \theta d\theta}{16^{5/2}(1-sin^2\theta)^{5/2}}\] = \[\int\frac{cos \theta d\theta}{256(cos^2\theta)^{5/2}}\]

OpenStudy (raspberryjam):

=\[\int\frac{cos \theta d\theta}{256(cos^5 \theta)}\] = \[\int\frac{d \theta}{256(cos^4 \theta)}\]

OpenStudy (raspberryjam):

\[\frac{1}{256}\int{sec^4\theta} d \theta\]

OpenStudy (raspberryjam):

Uhh oh man.. \[\frac{1}{256}\int{(1+tan^2\theta)sec^2\theta} d \theta\] = \[\frac{1}{256}\int{sec^2\theta+tan^2\theta sec^2\theta} d \theta\] then you have to split it =\[\frac{1}{256}\int{(tan^2\theta sec^2\theta) d\theta}+\int{sec^2\theta} d \theta\] use \[\theta=tan\theta\] and \[d\theta=sec^2\theta d\theta\] I think you should be able to get it from here.. It helps if you have this in front of you by the way: http://www.sosmath.com/trig/Trig5/trig5/trig5.html

OpenStudy (raspberryjam):

kind of skipped a step but you can see that \[\frac{1}{256}\int{sec^4\theta} d \theta\] = \[\frac{1}{256}\int{(sec^2\theta)(sec^2\theta)} d \theta\] and based on a trig identity \[sec^2\theta=1+tan^2\theta\]